Find the mass and centre of gravity of a triangular lamina with vertices (0,0), (1,0)and (0,2) if the density function is given by d(x, y) =1+ 3x + y.
1
Expert's answer
2019-03-05T11:51:44-0500

Solution

Equation of the hypotenuse of a right triangle have form

"y(x)=2-2x"


So mass of the triangular lamina

"M=\int\int\limits_{D}d(x,y)dxdy=\int\limits_{0}^{1}dx\int\limits_{0}^{2-2x}d(x,y)dy=""=\int\limits_{0}^{1}dx\int\limits_{0}^{2-2x}(1+ 3x + y)dy=\int\limits_{0}^{1}dx\frac12 (1+ 3x + y)^2\big|_0^{2-2x}=\int\limits_{0}^{1}dx\frac12 \big[(3-x)^2-(1+3x)^2\big]=""E=mc^2""=\int\limits_{0}^{1}dx(4-4x^2)=\frac83"

centre of gravity

"M_y=\frac1M\int\int\limits_{D}yd(x,y)dxdy=\frac1M\int\limits_0^{1}\frac16 y^2 (3 + 9 x + 2 y)\big|_{0}^{2-2x}="

"=\frac38\int\limits_0^{1}\frac23 (-1 + x)^2 (7 + 5 x)dx=\frac38\frac{11}{6}=\frac{11}{16}"

"M_x=\frac1M\int\int\limits_{D}xd(x,y)dxdy=\frac38\int\limits_{0}^{1}dx(4-4x^2)x=\frac38"

Answer: centre of gravity "(M_x,M_y)=(\frac38,\frac{11}{16})"

Mass: "M=\frac83"